OpenStudy (anonymous):

The function j(x) can only have a GCF factored out of it. then rewrite it as a group of factors if possible

OpenStudy (anonymous):

j(x) = x(x-3) the group of factors are x and x-3 as shown

OpenStudy (anonymous):

oh thank you so the original equation would be x^2 - 3x?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

can I ask you a couple other questions?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay so you dont have to help me with them all but here they areGiven the function k(x) = x2, compare and contrast how the application of a constant, c, affects the graph. The application of the constant must be discussed in the following manners: k(x + c) k(x) + c k(cx) c • k(x)

OpenStudy (anonymous):

k(x) +c just raises or lowers the graph c units; for example, y = x^2 - 5 will lower y= x^2 five units vertically; while y = x^2 + 2 will raise the parabola 2 units vertically. So, k(x) + c is a vertical shift, c units higher or lower, depending on the number c.

OpenStudy (anonymous):

k(x+c) will move the parabola y = x^2 horizontally c units. It will move it c units to the right if c<0; it will move it c units to the left when c>0. So this is a horizontal shift of c units.

OpenStudy (anonymous):

c k(x), the last one, is multiplying the function by c. If c>1, then it will narrow the original graph. while if 0<c<1, it will widen it. For example, y= 2x^2 is narrower than y = x^2, becuase we are multiplying x^2 by 2, so c = 2, and makes the graph narrower. While 1/2 x^2 will widen the parabola of y = x^2.

OpenStudy (anonymous):

thank you so much it is very helpful!

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Feel free to ask anything you dont understand. I tried to summarize it in a "short and sweet" form.

OpenStudy (anonymous):

i Appreciate it